### v1

@@ 4 Mechanism Choice
In this section, we show that another key result of Pathak and Sönmez (2008) no longer holds true in the multi-district setting. Specifically, we show that in the multi-district setting, a sophisticated student may strictly prefer for a district to use DA instead of BM. This is true regardless of whether the sophisticated student is constrained or unconstrained. In particular, we will give an example where the sophisticated student only finds schools in one district acceptable and prefers for that district to use DA.
4.1 Example: A sophisticated student may strictly prefer Deferred Acceptance
Our example is as follows. Suppose there are two districts with schools $\ell_{1},\ell_{2}\in L$ and school $r_{1}\in R$, where each school has unit capacity. Further suppose that there are three students $i_{1}$, $i_{2}$, and $i_{3}$.
The students’ preference orderings, districts of residence, constraint types, and sophistication types are as follows. Students $i_{1}$ and $i_{2}$ reside in $L$ and have arbitrary constraint types. Student $i_{3}$ is unconstrained and resides in an arbitrary district.
[…]
The schools’ priority orderings include the following:
@@ 4.1 Example: A sophisticated student may strictly prefer Deferred Acceptance
We will show that $i_{2}$ prefers district $L$ to use DA instead of BM. First, assume that district $L$ is using BM. As sincere student $i_{1}$ will not apply to $\ell_{2}$ in the first round, sophisticated student $i_{3}$ will choose to apply to $\ell_{2}$ in the first round, guaranteeing $i_{3}$’s acceptance at $\ell_{2}$. Sophisticated student $i_{2}$ will realize she has no chance at $\ell_{2}$ and will instead apply to $\ell_{1}$. This process results in $i_{1}$ unassigned, $i_{2}$ assigned to $\ell_{1}$, and $i_{3}$ assigned to $\ell_{2}$, which is the unique Nash equilibrium outcome. Note that in this Nash equilibrium outcome, $i_{2}$ is assigned to her second-choice school.
Suppose instead that district $L$ uses DA. Assume first that student $i_{2}$ uses her dominant strategy of ranking $\ell_{2}$ above $\ell_{1}$. This induces student $i_{3}$ to enroll in $R$ instead of $L$, as enrolling in $L$ would result in $i_{3}$ being unassigned. Specifically, $i_{3}$ could “knock out” $i_{2}$ from $\ell_{2}$, but in that case $i_{2}$ would knock out $i_{1}$ from $\ell_{1}$ and $i_{1}$ would in turn knock out $i_{3}$ from $\ell_{2}$. This constitutes a Nash equilibrium. Now consider the other possible strategies for $i_{2}$. If she ranks $\ell_{1}$ first, then she is assigned to $\ell_{1}$ for sure. This would lead to $i_{1}$ being matched to $\ell_{2}$ and hence $i_{3}$ enrolls in $R$ and is matched to $r_{1}$; in this case, $i_{2}$ has a profitable deviation to ranking $\ell_{2}$ first, so this is not a Nash equilibrium. Finally, if $i_{2}$ ranks only $\ell_{2}$, then it is easy to see that $i_{3}$ enrolls in $L$ and is matched to $\ell_{2}$, and $i_{2}$ becomes unassigned; in this case, $i_{2}$ has a profitable deviation of ranking $\ell_{1}$ somewhere in her preferences, so this is also not a Nash equilibrium. Altogether, the unique Nash equilibrium outcome has $i_{1}$ assigned to $\ell_{1}$, $i_{2}$ assigned to $\ell_{2}$, and $i_{3}$ assigned to $r_{1}$. In this Nash equilibrium outcome, $i_{2}$ is assigned to her first-choice school, which is a strict improvement for her compared to when district $L$ uses BM.
A key feature of this example is that there exists a cycle within the preferences of $i_{1}$ and $i_{2}$. This causes $i_{3}$ to get knocked out of $\ell_{2}$ when $i_{3}$ applies to district $L$ and district $L$ is using DA. We show in Appendix A that this example can be generalized to include a cycle with additional sophisticated students, in which each sophisticated student within the cycle similarly strictly prefers for her district to use DA.
@@ 4.2 Large-Market Analysis
We generalize the example from Section 4.1 to the uniform $(n;2)$ setting. (The same analysis also works in the uniform $(n;k)$ model for any constant $k\geq 2$.) We say that a sophisticated student $i$ prefers for a district $d$ to use DA if $i$ strictly prefers her match in every Nash equilibrium of the multi-district choice problem when $d$ uses DA to her match in every Nash equilibrium of the multi-district choice problem when $d$ uses BM. We show that there can be a constant number of sophisticated students who each prefer for the district that contains her entire preference list to use DA.
@@
For every $p\in(0,1)$, there exists $\tau>0$ such that for any large enough $n$, in the uniform $(n;2)$ model there exists a set of sophisticated students of expected size at least $\tau$ where each sophisticated student strictly prefers for the district that contains her entire preference list to use DA rather than BM, regardless of the mechanism used by the other district.
@@
Let the two districts be $L$ and $R$. We start by lower bounding the expected numbered of ordered triplets of students $(i_{1},i_{2},i_{3})$ that satisfy the following conditions (as in the example from Section 4.1):
@@
    Conditions on sophistication types, constraint types, and districts of residence:
@@
      $i_{1}$ most prefers a school $\ell_{1}\in L$, and next prefers a school $\ell_{2}\in L$.
@@
      $i_{3}$ most prefers $\ell_{2}$ and next prefers a school $r_{1}\in R$.
@@
      None of the other $n-3$ students finds $\ell_{1}$, $\ell_{2}$, or $r_{1}$ acceptable.
@@
As in the example in Section 4.1, in this situation, $i_{2}$ would strictly benefit from $L$ using DA rather than BM.
We want to determine the probability that this set of conditions occurs for a specific $(i_{1},i_{2},i_{3})$. Each of the three sets of conditions is independent. The conditions on sophistication types, constraint types, and districts of residence are satisfied with probability at least $p^{3}$. The conditions on student preferences and school locations are satisfied with probability
$\left(\frac{1}{2}\cdot\frac{1}{2}\right)\cdot\left(\frac{1}{n}\cdot\frac{1}{n-1}\right)\cdot\left(\frac{1}{n}\cdot\frac{n-2}{n-1}\cdot\frac{1}{2}\right)\cdot\left(\frac{n-3}{n}\cdot\frac{n-4}{n-1}\right)^{n-3}.$
In the expression above, each of the four parenthetical expressions corresponds to one of the four conditions on student preferences and school locations, and represents the condition’s probability, conditioned on the previous conditions. For sufficiently large $n$, the above expression can be lower bounded by
@@
The number of ordered triplets of distinct students is $n(n-1)(n-2)$. List all such triplets and let the $j$th element of the list be $I_{j}$. Let $\mathbbm{1}_{I_{j}}$ be the indicator random variable which represents whether the students in $I_{j}$ satisfy all of the conditions. Then the expected number of triplets of students that satisfy all conditions is
@@
To conclude the proof, we observe that by construction, no student can participate in two different triplets that satisfy the above conditions. Therefore, all such triplets of students involve distinct students, so there are at least an expected $\frac{p^{3}}{2^{6}\cdot 3\cdot e^{6}}$ sophisticated students who each strictly prefer for the district that contains their entire preference list to use DA over BM. We choose $\tau=\frac{p^{3}}{2^{6}\cdot 3\cdot e^{6}}$ to satisfy the theorem statement. ∎
@@ 5 Constraint Types
In the previous two sections, we have shown that two predictions made by Pathak and Sönmez (2008) no longer hold in multi-district settings. The first of these predictions revolved around whether a sophisticated student might prefer for another student to change their sophistication type in a particular way (specifically, becoming sophisticated). In our setting, students are not only characterized by their sophistication type but also by their constraint type, and hence it is natural to ask whether and when some students might prefer for some other students to change their constraint type.
We find a strong “anything goes” result here: For any combination of constraint types for two students, it might be the case that the first student strictly prefers for the constraint type of the second student to change, and this is furthermore abundant in large random markets. This holds regardless of the sophistication types of the two students, regardless of whether or not they reside in the same district, and regardless of the mechanisms used by the two districts. (We state this result in the uniform $(n;2)$ model; however, similarly to our previous results, the same analysis also works in the uniform $(n;k)$ model for any constant $k\geq 2$.) We say that a student $i_{a}$ prefers for student $i_{b}$ to change her constraint type if $i_{a}$ is strictly worse off in every Nash equilibrium of the multi-district school choice problem when $i_{b}$ has her given constraint type compared to every Nash equilibrium of the multi-district choice problem when $i_{b}$ has the opposite constraint type.
@@
For every $p\in(0,1)$, there exists $\tau>0$ such that for every pair of sophistication types $s_{1}$ and $s_{2}$, every pair of constraint types $c_{1}$ and $c_{2}$, and for any large enough $n$, in the uniform $(n;2)$ model with the districts using any combination of matching mechanisms, there exists a set of expected size at least $\tau n$ of students of sophistication type $s_{1}$ and constraint type $c_{1}$ where each student in the set strictly prefers for a distinct student of sophistication type $s_{2}$ and constraint type $c_{2}$ from the same district to change her constraint type. Furthermore, this also holds if “from the same district” is replaced with “from another district.”
@@
As it turns out, two types of constructions suffice to cover all of the various combinations of constraint types, sophistication types, districts of residence, and mechanisms in Theorem 5.1. Let $i_{1}$ and $i_{2}$ be two students, where $i_{1}$ is the student who prefers for $i_{2}$ to change her constraint type. Both constructions involve $i_{2}$ vacating her seat at a school $s$—either because she becomes constrained and $s$ is not in her district of residence, or because she becomes unconstrained and would rather enroll in another district.
@@
The simpler of the two constructions has $i_{1}$ filling the vacancy left by $i_{2}$ at $s$. This construction can be used in cases where $i_{1}$ is unconstrained, as well as in cases where $i_{1}$ is constrained and the school $s$ is in $i_{1}$’s district of residence. The second, slightly more elaborate construction covers the remaining cases, in which $i_{1}$ is constrained and the school $s$ is outside $i_{1}$’s district of residency. These cases, in which our results are arguably more surprising at first glance, include situations where $i_{2}$ resides in the same district as (the constrained) $i_{1}$ and vacates a spot in another district when $i_{2}$ becomes constrained, as well as situations where $i_{2}$ resides in a different district than $i_{1}$ and vacates a spot in that district when $i_{2}$ becomes unconstrained. In such cases, we introduce a third, unconstrained student, $i_{3}$, who takes the seat vacated by $i_{2}$ and therefore vacates a seat in the district of $i_{1}$, which $i_{1}$ in turn gets to fill.
@@
We now prove two cases of Theorem 5.1—one using each of the two constructions. The proof of each of the other cases of Theorem 5.1 is completely analogous to the proof of one of these two cases, as sketched above. We start by demonstrating the first, simpler construction.
Lemma 5.2.
For every $p\in(0,1)$, there exists $\tau>0$ such that for any large enough $n$, in the uniform $(n;2)$ model with the districts using any combination of matching mechanisms, there exists a set of unconstrained students of expected size at least $\tau n$ where each unconstrained student strictly prefers for a distinct constrained student to become unconstrained.

### v2

@@ 4 Mechanism Choice
In this section, we show that another key result of Pathak and Sönmez (2008) no longer holds true in the multi-district setting. Specifically, we show that in the multi-district setting, a sophisticated student may strictly prefer for her district to use DA instead of BM. This is true regardless of whether the sophisticated student is constrained or unconstrained. In particular, we give an example where the sophisticated student only finds schools in one district acceptable and prefers for that district to use DA.
4.1 Example: A sophisticated student may strictly prefer DA
Our example is as follows. Suppose there are two districts, $L$ and $R$, with schools $\ell_{1},\ell_{2}\in L$ and school $r_{1}\in R$ respectively, where each school has unit capacity. Further suppose that there are three students, $i_{1}$, $i_{2}$, and $i_{3}$.
The students’ sophistication types and preference orderings, as well as the schools’ priority orderings, are shown in the table below. Students $i_{1}$ and $i_{2}$ reside in $L$ and have arbitrary constraint types. Student $i_{3}$ is unconstrained and resides in an arbitrary district.
Student preferences
$\text{(sincere) }i_{1}:$ $\ell_{1}\succ\ell_{2}$
$\text{(sophisticated) }i_{2}:$ $\ell_{2}\succ\ell_{1}$
$\text{(sophisticated) }i_{3}:$ $\ell_{2}\succ r_{1}$
School priority orderings
@@ 4.1 Example: A sophisticated student may strictly prefer DA
We show that $i_{2}$ prefers for district $L$ to use DA rather than BM. First, assume that district $L$ is using BM. As sincere student $i_{1}$ does not apply to $\ell_{2}$ in the first round, sophisticated student $i_{3}$ chooses to apply to $\ell_{2}$ in the first round, guaranteeing $i_{3}$’s acceptance at $\ell_{2}$. Sophisticated student $i_{2}$ hence realizes that she has no chance at $\ell_{2}$ and instead applies to $\ell_{1}$. This process results in $i_{1}$ unassigned, $i_{2}$ assigned to $\ell_{1}$, and $i_{3}$ assigned to $\ell_{2}$, which is the unique Nash equilibrium outcome. Note that in this Nash equilibrium outcome, $i_{2}$ is assigned to her second-choice school.
Suppose instead that district $L$ uses DA. Then, student $i_{2}$ uses her dominant strategy of ranking $\ell_{2}$ above $\ell_{1}$. This induces student $i_{3}$ to enroll in $R$ instead of $L$, as enrolling in $L$ would result in $i_{3}$ being unassigned. Specifically, $i_{3}$ could “knock out” $i_{2}$ from $\ell_{2}$, but in that case $i_{2}$ would knock out $i_{1}$ from $\ell_{1}$ and $i_{1}$ would in turn knock out $i_{3}$ from $\ell_{2}$. This process results in $i_{1}$ assigned to $\ell_{1}$, $i_{2}$ assigned to $\ell_{2}$, and $i_{3}$ assigned to $r_{1}$, which is the unique Nash equilibrium outcome. In this Nash equilibrium outcome, $i_{2}$ is assigned to her first-choice school, which is a strict improvement for her compared to if district $L$ uses BM.
A key feature of this example is that there exists a cycle within the preferences of $i_{1}$ and $i_{2}$. This causes $i_{3}$ to get knocked out of $\ell_{2}$ when $i_{3}$ applies to district $L$ and district $L$ is using DA. We show in Appendix A that this example can be generalized to include a cycle with additional sophisticated students, in which each sophisticated student within the cycle similarly strictly prefers for the district containing her entire preference list to use DA.
@@ 4.2 Large-Market Analysis
We generalize the example from Section 4.1 to the uniform $(n;2)$ model. (The same analysis also works in the uniform $(n;k)$ model for any constant $k\geq 2$.) We say that a sophisticated student $i$ strictly prefers for a district $d$ to use DA if $i$ strictly prefers her match in every Nash equilibrium of the multi-district school choice problem when $d$ uses DA to her match in every Nash equilibrium of the multi-district school choice problem when $d$ uses BM. We show that there can be a constant number of sophisticated students who each prefer for the district that contains her entire preference list to use DA.
@@
For every $p\in(0,1)$, there exists $\tau>0$ such that for any large enough $n$, in the uniform $(n;2)$ model with category-probability lower bound $p$, there exists a set of sophisticated students of expected size at least $\tau$ wherein each sophisticated student strictly prefers for the district that contains her entire preference list to use DA rather than BM, regardless of the mechanism used by the other district.
@@
We start by lower bounding the expected numbered of ordered trios of students $(i_{1},i_{2},i_{3})$ that satisfy the following conditions (as in the example from Section 4.1):
@@
    Conditions on sophistication types, constraint types, and residence:
@@
      $i_{1}$ most prefers a school $\ell_{1}\in L$, and second most prefers a school $\ell_{2}\in L$.
@@
      $i_{3}$ most prefers $\ell_{2}$ and second most prefers a school $r_{1}\in R$.
@@
      No student finds a school in $\{\ell_{1},\ell_{2},r_{1}\}$ acceptable except as above.
@@
As in the example in Section 4.1, under these conditions, $i_{2}$ strictly benefits from $L$ using DA rather than BM.
We lower bound the probability that this set of conditions occurs for a specific $(i_{1},i_{2},i_{3})$. Each of the three sets of conditions is independent. The conditions on sophistication types, constraint types, and residence are satisfied with probability at least $p^{3}$. The conditions on student preferences and school locations are satisfied with probability at least
$\left(\frac{1}{2}\cdot\frac{1}{2}\right)\cdot\left(\frac{1}{n}\cdot\frac{1}{n-1}\right)\cdot\left(\frac{1}{n}\cdot\frac{n-2}{n-1}\cdot\frac{1}{2}\right)\cdot\left(\frac{n-3}{n}\cdot\frac{n-4}{n-1}\right)^{n}.$
In the expression above, each of the four parenthetical expressions corresponds to one of the four conditions on student preferences and school locations, and represents a lower bound on the condition’s probability, conditioned on the previous conditions. For sufficiently large $n$, the above expression can be lower bounded by
@@
The number of ordered trios of distinct students is n(n1)(n2)n(n-1)(n-2). List all such trios and let the $j$th element of the list be $I_{j}$. Let $\mathbbm{1}_{I_{j}}$ be the indicator random variable that represents whether the students in $I_{j}$ satisfy all of the above conditions. Then the expected number of trios of students that satisfy all conditions is
@@
We observe that by construction, it is impossible for there to be two different trios $(i_{1},i_{2},i_{3})$ and $(i^{\prime}_{1},i^{\prime}_{2},i^{\prime}_{3})$ that satisfy the above conditions such that $i^{\prime}_{2}=i_{2}$. Indeed, this would imply that $(i_{1},i_{2},i_{3})=(i^{\prime}_{1},i^{\prime}_{2},i^{\prime}_{3})$ because the identity of $i_{2}$ uniquely determines those of $i_{1}$ and $i_{3}$ (these are the only two other students who find acceptable the first choice of $i_{2}$, preferring it second and first respectively), while the identity of $i^{\prime}_{2}=i_{2}$ uniquely determines those of $i^{\prime}_{1}$ and $i^{\prime}_{3}$ in the same way.
To conclude, all such trios of students involve distinct $i_{2}$” students, so there are at least an expected $\frac{p^{3}}{2^{6}\cdot 3\cdot e^{6}}$ sophisticated students who each strictly prefer for the district that contains their entire preference list to use DA over BM. We choose $\tau=\frac{p^{3}}{2^{6}\cdot 3\cdot e^{6}}$ to satisfy the theorem statement. ∎
@@ 5 Constraint Types
In the previous two sections, we have shown that two predictions made by Pathak and Sönmez (2008) no longer hold in multi-district settings. The first of these predictions revolved around whether a sophisticated student might prefer for another student to change their sophistication type in a particular way (specifically, becoming sophisticated). In our setting, students are not only characterized by their sophistication type but also by their constraint type, and hence it is natural to ask, for completeness, whether and when some students might prefer for some other students to change their constraint type.
We find a strong “anything goes” result here: For any combination of constraint types for two students, it might be the case that the first student strictly prefers for the constraint type of the second student to change, and this is furthermore abundant in large random markets. This holds regardless of the sophistication types of the two students, regardless of whether or not they reside in the same district, and regardless of the mechanisms used by the two districts. (We state this result in the uniform $(n;2)$ model; however, similarly to our previous results, the same analysis also works in the uniform $(n;k)$ model for any constant $k\geq 2$.) We say that a student $i_{a}$ strictly prefers for a student $i_{b}$ to change her constraint type if $i_{a}$ is strictly worse off in every Nash equilibrium of the multi-district school choice problem when $i_{b}$ has her given constraint type compared to every Nash equilibrium of the multi-district school choice problem when $i_{b}$ has the opposite constraint type.
@@
For every $p\in(0,1)$, there exists $\tau>0$ such that for every pair of (not necessarily distinct) sophistication types $s_{1}$ and $s_{2}$, every pair of (not necessarily distinct) constraint types $c_{1}$ and $c_{2}$, and for any large enough $n$, in the uniform $(n;2)$ model with category-probability lower bound $p$ and with the districts using any combination of matching mechanisms, there exists a set of expected size at least $\tau n$ of students of sophistication type $s_{1}$ and constraint type $c_{1}$ wherein each student strictly prefers for a distinct student of sophistication type $s_{2}$ and constraint type $c_{2}$ from the same district to change her constraint type. Furthermore, this also holds if “from the same district” is replaced with “from the other district.”
@@
As it turns out, two types of constructions suffice to cover all of the various combinations of constraint types, sophistication types, districts of residence, and mechanisms in Theorem 5.1. Let $i_{1}$ and $i_{2}$ be two students, where $i_{1}$ is the student who prefers for $i_{2}$ to change her constraint type. Both constructions involve $i_{2}$ vacating her seat at a school ss—either because she becomes constrained and $s$ is not in her district of residence, or because she becomes unconstrained and would rather enroll in another district.
@@
The simpler of the two constructions has $i_{1}$ filling the vacancy left by $i_{2}$ at $s$. This construction can be used in cases where $i_{1}$ is unconstrained, as well as in cases where $i_{1}$ is constrained and the school $s$ is in $i_{1}$’s district of residence. The second, slightly more elaborate construction covers the remaining cases, in which $i_{1}$ is constrained and the school $s$ is outside $i_{1}$’s district of residence. These cases, in which our results are arguably more surprising at first glance, include situations where $i_{2}$ resides in the same district as (the constrained) $i_{1}$ and vacates a spot in another district when $i_{2}$ becomes constrained, as well as situations where $i_{2}$ resides in a different district than $i_{1}$ and vacates a spot in that district when $i_{2}$ becomes unconstrained. In such cases, we introduce a third, unconstrained student, $i_{3}$, who takes the seat vacated by $i_{2}$ and therefore vacates a seat in the district of $i_{1}$, which $i_{1}$ in turn gets to fill.
@@
In Appendix B, we prove two cases of Theorem 5.1—one using each of the two constructions. The proof of each of the other cases of Theorem 5.1 is completely analogous to the proof of one of these two cases.
6 Sincere-Unconstrained Students
In this section, we dive deeper into the machinery behind the results of Section 3, and ask which features of the school choice problems used in the analysis of that Section are necessary to give rise to the phenomenon of sophisticated students who prefer for sincere students to become sophisticated.
The definitions of three of the student types in our taxonomy follow closely from ideas in Pathak and Sönmez (2008): the two constrained types behave, within their district, just like the two types in Pathak and Sönmez (2008), while sophisticated-unconstrained students have complete strategic freedom, analogous to sophisticated students having complete strategic freedom within the framework of Pathak and Sönmez (2008). Our fourth type—sincere-unconstrained students—while still heavily inspired by Pathak and Sönmez (2008) and completely in line with our taxonomy, is conceptually (and formally, as we discuss below) somewhat further from the types defined by that paper. Therefore, we first ask whether and to what extent this student type plays a key role in our analysis, and then ask whether and to what extent our results are robust to changes in the precise definition of such students.
6.1 The Role of Sincere-Unconstrained Students in the Results of Section 3
6.1.1 Example: When all sincere students are constrained
The construction in Section 3 hinges on a trait that only a sincere-unconstrained student can possess: She can block a sophisticated student from getting a seat at some school, and yet, she may vacate that seat for another school if she becomes sophisticated. [^16] While a sincere-constrained student can also block a sophisticated student in the same way, a sincere-constrained student will not vacate that seat if she becomes sophisticated, as it is the best outcome she can achieve. The first question we ask is whether there exists a school choice problem in which a sophisticated student prefers for a sincere student to become sophisticated even in the absence of sincere-unconstrained students. We give an affirmative answer, whose construction is the most involved in this paper so far.
Proposition 6.1.
There exists a two-district school choice problem in which (1) all sincere students are constrained, and (2) there exists a sophisticated student who strictly prefers for a specific sincere student to become sophisticated.